# Explicit Multimodal Graph Modeling for Human-Object Interaction Detection

### v1

@@ Introduction
As illustrated in Figure 2, the process begins with an off-the-shelf object detector, which extracts bounding boxes and appearance features. Based on these detection results and a matching strategy, we construct human and object instances $\mathbf{N}$, along with human features $\mathbf{H}$ and object features $\mathbf{O}$. At the core of our framework lies the MFI mechanism, as detailed in Algorithm 1 (Lines 1–7). The process commences by constructing the initial human-object pair features, which are then iteratively refined through the proposed four-stage pipeline. Following (Park, Park, and Lee 2023), we set the number of iterative steps $T$ to 2. Notably, in the visual and textual stages, we leverage visual and textual features to update the human and object instances, thereby infusing multimodal information into the node encodings. In addition, since the message passing is guided by a weight matrix computed from the relational structure among human-object pairs, this implicitly promotes feature interaction between pairs. In contrast, the spatial and interaction stages are explicitly designed to enhance feature communication between human-object pairs. Finally, a decoder is employed to incorporate contextual semantic information and produce the final predictions. In this decoder, the human-object pair features act as queries, while the global features extracted from the backbone serve as keys and values.
Experiments
@@
Full Rare Non-Rare Full Rare Non-Rare $AP^{S1}_{role}$ $AP^{S2}_{role}$
@@
UnionDet (2020) R50-F 17.58 11.72 19.33 19.76 14.68 21.27 47.5 56.2
PPDM (2024) Swin-B-F 27.59 18.07 30.44 29.11 18.85 32.71 - -
AS-Net (2021) R50 28.87 24.25 30.25 31.74 27.07 33.14 53.9 -
FGAHOI-S (2024b) Swin-T 29.94 22.24 32.24 32.48 24.16 34.97 60.5 61.2
DOQ (2022) R50+C 33.28 29.19 34.50 - - - 63.5 -
GEN-VLTK (2022) R50+C 33.75 29.25 35.10 36.78 32.75 37.99 62.4 64.4
DiffHOI-S (2023) R50+C+SD 34.41 31.07 35.40 37.31 34.56 38.14 61.1 63.5
FGAHOI-L (2024b) Swin-L 37.18 30.71 39.11 38.93 31.93 41.02 - -
DiffHOI-L (2023) Swin-L+C+SD 40.63 38.10 41.38 43.14 40.24 44.01 65.7 68.2
@@
SCG (2021) R50-F 29.26 24.61 30.65 32.87 27.89 34.35 54.2 60.9
PPDM++ (2024) Swin-B-F 30.10 23.73 32.00 31.80 24.93 33.85 - -
UPT (2022) R101 32.31 28.55 33.44 35.65 31.60 36.86 60.7 66.2
ADA-CM (2023) R50+C 33.80 31.72 34.42 - - - 60.7 66.2
ILCN (2025) Swin-T 33.80 29.83 34.99 - - - 63.4 65.3
PViC w/ DETR (2023) R50 34.69 32.14 35.45 38.14 35.38 38.97 62.8 67.8
HOIGen (2024) R50+C 34.84 34.52 34.94 - - - - -
ViPLO (2023) R101-D+C+VP 34.95 33.83 35.28 38.15 36.77 38.56 60.9 66.6
TKCE (2025) R50+C 35.11 34.35 35.34 - - - - -
LAIN (2025) R50+C 36.02 35.70 36.11 - - - 63.4 65.3
MP-HOI-S (2024) R50+C+SD 36.50 35.48 36.80 - - - 66.2 67.6
PViC w/ $\mathcal{H}$-DETR (2023) Swin-L 44.32 44.61 44.24 47.81 48.38 47.64 64.1 70.2
MP-HOI-L (2024) Swin-L+C+SD 44.53 44.48 44.55 - - - - -
MGNM R50+C 36.68 34.10 37.45 39.99 37.15 40.84 62.1 67.5
MGNM w/ $\mathcal{H}$-DETR Swin-L+C 46.54 48.09 46.07 49.78 50.82 49.47 64.3 70.5
Table 1: Performance comparison in terms of mAP (%) on the HICO-DET and V-COCO dataset. To ensure a fair comparison, we report both the performance and the corresponding feature backbone for each method. In each evaluation metric, the best result is marked in bold and the second-best result is underlined. In this paper, R50, R101, R50-F and R101-D denotes ResNet50, ResNet101, ResNe50-FPN and ResNet101-DC5 model, respectively. A similar notation applies to Swin Transformer (Liu et al. 2021). C, SD, and VP mean CLIP, Stable Diffusion, and ViTPose models, respectively. We note that MP-HOI (Yang et al. 2024) utilizes additional training data (Magic-HOI and SynHOI), which should be considered when comparing results.
@@
This section presents a series of experiments designed to evaluate the effectiveness of our proposed method. First, we briefly introduce the experimental settings. Next, we present a comprehensive comparison of our method against existing state-of-the-art methods. Subsequently, the ablation study of our method is conducted. Finally, we rethink and analyze the Rare-Non-rare bias in the HOI detection field and provide some promising suggestions.
Experimental Settings
Datasets
Our experiments are conducted on two widely used public benchmarks: HICO-DET and V-COCO. HICO-DET contains 38,118 training images and 9,658 test images. It includes 80 object classes from the MS COCO dataset (Lin et al. 2014) and 117 action classes, which combine to form 600 types of HOI instances. V-COCO, a smaller-scale subset of the MS COCO dataset, contains 2,533 training, 2,867 validation, and 4,946 test images, encompassing 80 object classes and 29 action classes. Given its larger scale and closer alignment with real-world data distributions, our primary experimental focus is on the HICO-DET dataset.
Evaluation Metrics
Following the standard evaluation protocols in (Park, Park, and Lee 2023), we report the mean Average Precision (mAP) for evaluation. An HOI triplet is considered a true positive when the IoU of the detected human and object is larger than 0.5 and the action is predicted accurately. For the HICO-DET dataset, we report performance under two standard evaluation settings: Default and Known Object. Each setting has three different subsets: (1) all 600 HOI classes (Full), (2) 138 HOI classes with less than 10 training instances (Rare), and (3) 462 HOI classes with 10 or more training instances (Non-Rare). For the V-COCO dataset, we evaluate performance based on two scenarios: Scenario 1, which requires correct prediction of occluded object bounding boxes, and Scenario 2, which does not.
@@
For HICO-DET and V-COCO datasets, we use the fine-tuned DETR and $\mathcal{H}$-DETR (Jia et al. 2023) weights provided by Zhang et al. (Zhang et al. 2023). During training, following the commonly used training paradigm of two-stage methods, the weights of the object detector are frozen, while only the interaction predictor is trained. Meanwhile, the focal loss (Lin et al. 2017) is employed as the loss function. Following (Lei et al. 2023), we also incorporate the adapter into the CLIP module. The model is optimized using the AdamW optimizer with an initial learning rate of $10^{-4}$. All models are trained for 30 epochs on 4 NVIDIA A100 GPUs, using a total batch size of 32. Unless otherwise specified, all other hyperparameters adhere to the settings established in (Zhang et al. 2023; Lei et al. 2023).
Comparison with State-of-the-art Methods
@@
Table 1 provides a comprehensive performance comparison between our proposed MGNM and current state-of-the-art methods. For clarity, the competing methods are grouped into one-stage and two-stage categories. On the HICO-DET dataset, MGNM outperforms all other one-stage and two-stage approaches that utilize a ResNet backbone. Specifically, in the Default Full setting, our method achieves a 2.27 mAP improvement over DiffHOI-S, despite the latter leveraging two VLMs, namely CLIP and Stable Diffusion (Rombach et al. 2021). MGNM also surpasses the leading two-stage approach, MP-HOI-S, by 0.18 mAP in the Default Full setting, even though MP-HOI-S incorporates both additional VLMs and extra training data. Moreover, MGNM exceeds the performance of the recent GNN-based ViPLO by 1.73 mAP in the Default Full setting, notwithstanding ViPLO’s adoption of the supplementary ViTPose-L model (Xu et al. 2022). In the Known Object setting, MGNM consistently outperforms all ResNet-based counterparts across all three subsets by a substantial margin. Notably, our ResNet50-based MGNM even surpasses FGAHOI-L, which utilizes a more powerful Swin-L backbone, by 1.06 mAP in the Full setting. Collectively, these results demonstrate the effectiveness of our multimodal graph network and its explicit modeling strategy for HOI detection. On the V-COCO dataset, MGNM also achieves highly competitive performance among ResNet-based methods. We note that MP-HOI-S obtains a higher $AP_{role}^{S1}$, which we attribute to its use of additional training data. This observation suggests that GNN-based approaches, including MGNM, could further benefit from large-scale pre-training, highlighting a promising direction for future research.
@@
w/o Spatial Stage 36.00 34.68 36.39
w/o Visual Stage 35.32 32.66 36.12
w/o Textual Stage 36.16 33.81 36.86
w/o Interaction Stage 35.61 33.37 36.28
vanilla 36.68 34.10 37.45
Table 2: Effectiveness of each component within our MFI mechanism. w/o means “without”.
Furthermore, to further demonstrate the scalability of our approach, we compare the performance of MGNM equipped with a more powerful object detector against other state-of-the-art methods that utilize similarly strong backbones. As shown in Table 1, MGNM integrated with $\mathcal{H}$-DETR achieves substantial performance improvements over the competing methods. For example, on the HICO-DET Default Full setting, our method outperforms PViC with $\mathcal{H}$-DETR by 2.22 mAP. Notably, although both DiffHOI-L and MP-HOI-L leverage two VLMs each, our single-VLM approach achieves significant gains of 5.91 and 2.01 mAP over these methods, respectively. These results underscore not only the scalability of our framework but also the effectiveness of our multimodal graph network design.
Ablation Study
@@
(b) ride snowboard
(c) drive boat
(d) Visualization for c
(e) Visualization for c
Figure 3: Qualitative results on the HICO-DET test set. Figures 3(a) and 3(b) illustrate successful predictions across a variety of challenging scenarios. In contrast, Figure 3(c) presents a representative failure case, wherein the model generates a false positive triplet. Specifically, it incorrectly infers that both individuals are driving the boat, whereas only the person on the right is performing this action. To further investigate this failure case, Figures 3(d) and 3(e) visualize the corresponding attention maps, highlighting the model’s regions of focus that led to this prediction.
Methods Backbone mAP $\uparrow$ Time $\downarrow$ Memory $\downarrow$
SCG R101 29.26 182 ms 7,702 MiB
ViPLO R101-D+C+VP 34.95 166 ms 10,224 MiB
MGNM R50+C 36.68 125 ms 2,946 MiB
Table 3: Comparison of mAP, inference time, and GPU memory on the HICO-DET test set. We select two prominent GNN-based baselines for this comparison: SCG and ViPLO. It is important to note that the official implementations of both SCG and ViPLO execute object detection and interaction prediction as two separate, sequential processes. Therefore, to ensure a fair comparison, our reported time and memory for these baselines encompass the complete process. All experiments are conducted with a batch size of 1 on a single NVIDIA A100 GPU. Time refers to the elapsed time required for processing one image.
In this subsection, we conduct ablation studies to analyse the effectiveness of our proposed components and the efficiency of our MGNM framework. All the experiments are conducted on the HICO-DET test set under Deafult setting.
@@
We first analyze the contribution of each of the four stages within our MFI mechanism, which collectively process low-level and high-level multimodal features. As shown in Table 2, removing the low-level feature stages, namely the Spatial (row 1) and Textual (row 3) stages, results in performance drops of 0.68 and 0.52 mAP, respectively. This demonstrates that these low-level spatial and textual features provide a valuable supplement to the foundational human-object pair representations. Ablating the high-level feature stages, the Visual (row 2) and Interaction (row 4) stages, induces a more substantial performance degradation, with drops of 1.36 and 1.07 mAP, respectively. These decreases are considerably larger than those observed in low-level stages. This finding suggests that the high-level features are more critical to our model’s predictive power, an observation that aligns with the notion that GNNs are inherently biased towards modeling high-level structural information (Antoun and Asmar 2023). Interestingly, we observe a counter-intuitive phenomenon in row 1: ablating the spatial stage improves performance on the rare classes. This suggests that low-level vision features may inadvertently amplify the model’s existing bias towards more frequent, non-rare classes. Nevertheless, considering their substantial positive impact on the Full and Non-Rare settings, these spatial features remain an indispensable component for overall model performance.
Efficiency Analysis
This subsection presents an efficiency analysis of MGNM in comparison to two related GNN-based methods, focusing on inference time and GPU memory usage. As reported in Table 3, both SCG and ViPLO decouple the object detection and interaction prediction processes, whereas our method integrates these stages into a unified end-to-end pipeline. To ensure a fair comparison, we report the total inference time and GPU memory for SCG and ViPLO, excluding file I/O operations. As shown in Table 3, our approach not only achieves substantially higher mAP but also demonstrates superior efficiency, recording the lowest inference time and GPU memory consumption among the compared methods. These results highlight the strong balance between effectiveness and efficiency achieved by MGNM. It is worth noting that the official object detection scripts provided by the baseline methods appear to be unoptimized with respect to memory usage. When replacing their detection modules with DETR weights, resulting in less GPU consumption than the actual GPU memory usage, the measured occupancy drops to 1,904 MiB for SCG and 4,426 MiB for ViPLO. Even under these conditions, when comprehensively considering both mAP and inference time, our method continues to deliver the best overall effectiveness and efficiency.
@@
Rethinking Rare-Non-rare Bias

### v2

@@ 3.2 Multimodal Graph Network Modeling
As illustrated in Figure 2, the process begins with an off-the-shelf object detector, which extracts bounding boxes and appearance features. Based on these detection results and a matching strategy, we construct human and object instances $\mathbf{N}$, along with human features $\mathbf{H}$ and object features $\mathbf{O}$. At the core of our framework lies the MFI mechanism, as detailed in Algorithm 1 (Lines 1–7). The process commences by constructing the initial human-object pair representations, which are then iteratively refined through the proposed four-stage pipeline. Following [park2023viplo], we set the number of iterative steps $T$ to 2. Notably, in the visual and textual stages, we leverage visual and textual features to update the human and object instances, thereby infusing multimodal information into the node encodings. In addition, since the message passing is guided by a weight matrix computed from the relational structure among human-object pairs, these two stages implicitly promote feature interaction between pairs. In contrast, the spatial and interaction stages are explicitly designed to enhance feature communication between human-object pairs. Finally, a decoder is employed to incorporate contextual semantic information and produce the final predictions. In this decoder, the human-object pair features act as queries, while the global features extracted from the backbone serve as keys and values.
4 Experiments
@@
Full Rare Non-Rare Full Rare Non-Rare $\text{AP}^{S1}_{role}$ $\text{AP}^{S2}_{role}$
@@
UnionDet [kim2020uniondet] R50 17.58 11.72 19.33 19.76 14.68 21.27 47.5 56.2
AS-Net [chen_2021_asnet] R50 28.87 24.25 30.25 31.74 27.07 33.14 53.9 -
FGAHOI [10315071] Swin-T 29.94 22.24 32.24 32.48 24.16 34.97 60.5 61.2
DOQ [qu2022distillation] R50+CLIP 33.28 29.19 34.50 - - - 63.5 -
DiffHOI [yang2023boosting] R50+CLIP+SD 34.41 31.07 35.40 37.31 34.56 38.14 61.1 63.5
MP-HOI [Yang2024OpenWorldHI] R50+CLIP+SD 36.50 35.48 36.80 - - - 66.2 67.6
FGAHOI-L [10315071] Swin-L 37.18 30.71 39.11 38.93 31.93 41.02 - -
DiffHOI-L [yang2023boosting] Swin-L+CLIP+SD 40.63 38.10 41.38 43.14 40.24 44.01 65.7 68.2
$\text{MP-HOI-L}^{\dagger}$ [Yang2024OpenWorldHI] Swin-L+CLIP+SD 44.53 44.48 44.55 - - - - -
@@
SCG [zhang2021scg] R50 29.26 24.61 30.65 32.87 27.89 34.35 54.2 60.9
PPDM++ [10496247] Swin-B 30.10 23.73 32.00 31.80 24.93 33.85 - -
UPT [zhang2022upt] R101 32.31 28.55 33.44 35.65 31.60 36.86 60.7 66.2
ILCN [lu2025intra] Swin-T 33.80 29.83 34.99 - - - 63.4 65.3
PViC [zhang2023pvic] R50 34.69 32.14 35.45 38.14 35.38 38.97 62.8 67.8
HOIGen [Guo2024UnseenNM] R50+CLIP 34.84 34.52 34.94 - - - - -
ViPLO [park2023viplo] R101-DC+CLIP+ViTPose 34.95 33.83 35.28 38.15 36.77 38.56 60.9 66.6
TKCE [10889833] R50+CLIP 35.11 34.35 35.34 - - - - -
LAIN [kim2025locality] R50+CLIP 36.02 35.70 36.11 - - - 63.4 65.3
VRDiff [Cao_2025_ICCV] R50+CLIP+SD 36.77 35.66 37.11 39.98 38.69 40.36 - -
$\text{ADA-CM}^{\dagger}$ [ting2023hoi] R50+CLIP 38.40 37.52 38.66 - - - 58.6 64.0
RLIPv2 [Yuan2023RLIPv2FS] Swin-T+BLIP 38.60 33.66 40.07 - - - 68.8 70.8
HORP [11092421] R50+CLIP+Gaze 38.61 36.14 39.34 40.98 38.25 41.79 65.6 68.3
$\text{HOLa}^{\dagger}$ [lei2025hola] R50+CLIP+GPT 39.05 38.66 39.17 - - - - -
PViC-L [zhang2023pvic] Swin-L 44.32 44.61 44.24 47.81 48.38 47.64 64.1 70.2
Pose-aware [wu2024exploring] Swin-L+ViTPose 46.01 46.74 45.80 49.50 50.59 49.18 63.0 68.7
HORP-L [11092421] Swin-L+CLIP+Gaze 47.53 46.81 47.74 51.24 50.78 51.38 68.9 71.1
MGNM R50+CLIP 39.25 38.66 39.42 42.36 41.76 42.54 67.7 70.2
$\text{MGNM}_{L}^{\dagger}$ Swin-L+CLIP 49.09 50.62 48.63 52.78 53.89 52.45 70.4 73.9
Table 1: Performance comparison in terms of mAP (%) on the HICO-DET and V-COCO dataset. In each evaluation metric, the best result is marked in bold and the second-best result is underlined. In this paper, R50, R101, and R101-DC denotes ResNet50, ResNet101, and ResNet101 with dilated convolution, respectively. SD means Stable Diffusion model [Rombach2021HighResolutionIS]. $\dagger$ means the CLIP adopts the ViT-L image encoder [Dosovitskiy2020AnII]. Others adopt the ViT-B image encoder.
@@
This section presents a series of experiments designed to evaluate the effectiveness of our proposed method. First, we briefly introduce the experimental settings. Next, we present a comprehensive comparison of our method against the SOTA methods. Subsequently, we conduct the ablation study of our method. All of its experiments are conducted on the HICO-DET dataset under Deafult setting. Finally, we analyze the Rare-Non-rare bias in the HOI detection field and provide some promising suggestions.
4.1 Experimental Settings
Datasets and Evaluation Metrics
Our experiments are conducted on two widely used public benchmarks: HICO-DET and V-COCO. Following the standard evaluation protocols in [zhang2023pvic, park2023viplo], we report the mean average precision (mAP) for evaluation. Given its larger scale and closer alignment with real-world data distributions, our primary experimental focus is on the HICO-DET dataset. More details about datasets and evaluation metrics can be found in the supplementary materials.
@@ 4.1 Experimental Settings
For HICO-DET and V-COCO datasets, following [zhang2023pvic, 11092421], we use the fine-tuned DETR and $\mathcal{H}$-DETR [Jia2022DETRsWH] models as the object detector. During training, following the commonly used training paradigm of two-stage methods, the weights of the object detector are frozen, while only the interaction predictor is trained. Meanwhile, the focal loss [Lin2017FocalLF] is employed as the loss function. Following [ting2023hoi, lei2025hola], we employ the CLIP ViT-B/16 and ViT-L/14@336px models for our method, and all the parameters of them are frozen, except for the adapter layer. We also apply data augmentation techniques from previous works [ting2023hoi, zhang2023pvic], including random cropping, resizing, and color jittering. During inference, following [zhang2023pvic, zhang2022upt], we use the hyparameter $\lambda$ to suppress overconfident objects. The model is optimized using the AdamW optimizer with an initial learning rate of $10^{-4}$. All models are trained for 30 epochs on 4 NVIDIA A100 GPUs, using a batch size of 16.
4.2 Comparison with State-of-the-art Methods
@@
Table 1 provides a comprehensive performance comparison between our proposed MGNM and current SOTA methods. For clarity, the competing methods are grouped into one-stage and two-stage categories. On the HICO-DET dataset, MGNM outperforms all other one-stage and two-stage approaches that utilize a ResNet backbone. Specifically, in the Default Full setting, our method achieves a 4.84 and 2.75 mAP improvement over the Transformer-based methods DiffHOI and MP-HOI, despite they leveraging two VLMs, namely CLIP and Stable Diffusion. MGNM also surpasses the leading two-stage approaches, HORP and HOLa, by 0.64 and 0.2 mAP in the Default Full setting, even though they incorporates additional models. Moreover, MGNM exceeds the performance of the recent GNN-based method, ViPLO, by 4.3 mAP in the Default Full setting, notwithstanding ViPLO’s adoption of the supplementary ViTPose-L model [xu2022vitpose]. In the Known Object setting, MGNM consistently outperforms all ResNet-based counterparts across all three subsets by a substantial margin. Notably, our R50-based MGNM even surpasses FGAHOI-L, which utilizes a more powerful Swin-L backbone [Liu2021SwinTH], by 3.43 mAP in the Full setting. Collectively, these results demonstrate the effectiveness of our multimodal graph network and its explicit modeling strategy for HOI detection. On the V-COCO dataset, MGNM also achieves substantial performance improvement among ResNet-based methods. We note that RLIPv2 obtains a higher performance on V-COCO, which we attribute to its large-scale pre-training on additional data. This observation suggests that GNN-based approaches, including MGNM, could further benefit from large-scale pre-training, highlighting a promising direction for future research.
To further demonstrate the scalability of our approach, we compare the performance of MGNM equipped with a more advanced object detector against other SOTA methods that utilize similarly strong backbones. As shown in Table 1, $\text{MGNM}_{L}$ achieves substantial performance improvements over the competing methods. For example, on the HICO-DET Default Full setting, our method outperforms the Transformer-based methods, PViC-L and Pose-aware, by 4.77 and 3.08 mAP. Notably, although both DiffHOI-L and MP-HOI-L leverage two VLMs each, our single-VLM approach achieves significant gains of 8.46 and 4.56 mAP over these methods, respectively. These results underscore not only the scalability of our framework but also the effectiveness of our multimodal graph network design.
@@
w/o Spatial Stage 38.17 37.62 38.33
w/o Visual Stage 36.78 35.64 37.12
w/o Textual Stage 38.23 37.44 38.47
w/o Interaction Stage 36.94 35.18 37.47
Simple Prompt 37.53 35.92 38.01
MLP Fusion 37.08 35.16 37.65
vanilla 39.25 38.66 39.42
Table 2: Effectiveness of each component within our framework. w/o means “without”.
4.3 Ablation Study
@@
(b) carry backpack
(c) sit on bench
(d) wear tie
(e) ride snowboard
(f) drive boat
(g) Visualization for f
(h) Visualization for f
Figure 3: Qualitative results on the HICO-DET dataset. Figures 3(a)-3(e) illustrate successful predictions across a variety of challenging scenarios. Figure 3(f) presents a representative failure case, and Figures 3(g) and 3(h) visualize its corresponding attention maps.
Methods Full Rare Non-Rare
CLIP-ViT-B/32 38.83 38.24 39.01
CLIP-ViT-B/16 39.25 38.66 39.42
CLIP-ViT-L/14 42.97 42.32 43.17
CLIP-ViT-L/14@336px 43.53 42.88 43.73
Table 3: Comparison of our method with different CLIP sizes.
@@
We first analyze the contribution of each of the four stages within our MFI mechanism, which collectively process low-level and high-level multimodal features. As shown in Table 2, in row 1 (w/o Spatial Stage), we only ablate the spatial features, preserving the 3D location prior and weight matrix. This results in performance drops of 1.08 mAP. In row 3 (w/o Textual Stage), it leads to a performance drop of 1.02 mAP. This phenomenon demonstrates that these low-level spatial and textual features provide a valuable supplement to the foundational human-object pair representations. Ablating the high-level feature stages, the Visual (row 2) and Interaction (row 4) stages, induces a more substantial performance degradation, with drops of 2.47 and 2.31 mAP, respectively. These decreases are considerably larger than those observed in low-level stages. This finding suggests that the high-level features are more critical to our model’s predictive power, an observation that aligns with the notion that GNNs are inherently biased towards modeling high-level structural information [antoun2023human]. Next, we analyze the other two important components in our method. In row 5 (Simple Prompt), to analyze the effectiveness of the interaction-centric prompt, we replace the interaction-centric prompt with a simple prompt “a photo of person interacting with $\langle\text{object}\rangle$”. As shown in Table 2, it results in a performance drop of 1.72 mAP. This demonstrates the effectiveness of our interaction-centric prompt, which provides more fine-grained semantic cues for the text encoder. At last, in row 6 (MLP Fusion), we ablate the GNN structure in our method and only use a simple MLP fusion to fuse the corresponding multimodal features for the instances. This can be interpreted as a non-GNN baseline. It can be observed that this ablation results in a drop of 2.17 mAP. This demonstrates the significance of our explicit modeling strategy. Based on the above analysis, the systematic ablation studies collectively demonstrates the effectiveness of our explicit multimodal graph network modeling.
Analysis on CLIP Size
CLIP, as a classic vision language model, has been widely adopted in HOI detection and demonstrated an effective model [ting2023hoi, Lei2024EZHOIVA, park2023viplo]. Since different versions of CLIP differ significantly in model architecture and parameter count, in this subsection, we analyze the influence of different CLIP model scales on our method. As shown in Table 3, we find that CLIP with the same architecture achieves competitive performance. For instance, CLIP-ViT-B/16 outperforms CLIP-ViT-B/32 by 0.42 mAP in the Full setting. This small margin is expected. While ViT-B/16 produces more detailed visual features, they share the same base architecture, and the parameter count is almost same. Furthermore, other modules in our HOI detector also provide fine-grained features, which likely further mitigates the gap between the ViT-B/32 and ViT-B/16 variants. In contrast, a significant difference emerges between the ViT-B and ViT-L model families. For example, CLIP-ViT-L/14 achieves a significant improvement of 3.72 mAP over CLIP-ViT-B/16. This is expected, as ViT-L/14 has a larger architecture and possesses approximately $3.5\times$ as many parameters as ViT-B/16, resulting in a much larger model capacity. Additionally, the text encoders in different CLIP sizes are architecturally similar, varying primarily in their projection dimension. Consequently, the improvement is primarily attributable to the more powerful visual encoder. Based on this analysis, adopting a larger CLIP model appears to be an effective strategy for improving performance in HOI detection. We also conduct ablation studies on CLIP sizes with the $\mathcal{H}$-DETR. Although the performance gain is not as significant as that with R50, it still brings an effective improvement of 1.17 mAP. More discussion about this can be found in the supplementary materials.
@@
4.4 Rare-Non-rare Bias